At 88 meters high a rock is thrown horizontally from the top of a cliff, with a horizontal speed of 25 m/s. Now, what's the direction of velocity and with what velocity does the rock hit the ground?

1
Expert's answer
2022-06-08T17:25:45-0400

"v_x=25~\frac ms,"

"v_y=\sqrt{2gh}=41.5~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=48.5~\frac ms,"

"\varphi=-\arctan\frac{v_y}{v_x}=-59°."